1. In an office with 21 staff members, 1/3 are men and 2/3 are women. To obtain a staff in which 1/4 are  men, how many women should be hired? 

Grade:10

4 Answers

soumya ranjan prusty
5 years ago
03...............................................................................................................................................................................................................................................................................................................................................................................
Darpan Verma
5 years ago
If in a staff 21 members are there then 7 will be men and 14 will be women.Let the no.of staff after increasing be y.so we will have y/4 = 7 because the no.of men are fixed . By solving we find y = 28.From which 7 are men so women will be 21.earlier womens are 14.Therefore the increased no. is 7.
Arun
5 years ago
The given info tells you there are 21 employees and 1/3 are men. 
21 * 1/3 = 7
so there are 7 men. Thus there are 14 women 
You are asked to find out ho wmany women to hire such that men wind up being 1/4 of the total 
 

The number of men won't change. 
7*4 = 28
We need to get a total of 28 employees.
Since there are 21 now, and we need 28, hire 7 women.
Nitesh
5 years ago
Answer is 7 more woman required.Total staff=21no. of men =21×1/3=7no. of woman =21×2/3=14Let some more woman hired is X so that no. of man becomes 1/4 and no. of woman becomes 3/4Now the total no of staff becomes 21+XIt is given that no. of man/ total staff=1/4So 7/21+X=1/4=》21+X=28=》X=7So 7 more woman hired.